OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve f over g(−3) and select the correct answer below: f(x) = 6x + 8 g(x) = x − 2 −2 −negative one half one half 2

OpenStudy (anonymous):

1 Find the Greatest Common Factor (GCF) 1. What is the largest number that divides evenly into 6x and 8? 2 2. What is the highest degree of x that divides evenly into 6x and 8? None, x is not in every term 3. Multiplying the results above, the GCF is 2 GCF=2 2 Factor out the GCF 1. Put the GCF as the first term 2. Then, in parentheses, divide each term by the GCF 2(6x2+82) 3 Simplify each term in parentheses 2(3x+4)

OpenStudy (anonymous):

\[\frac{ f }{g }(x)=\frac{ f(x) }{ g(x)}=\frac{ 6x+8 }{ x-2}\] Plug in -3 for x

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i did 6(-3)+8 ----------- -3-2

OpenStudy (anonymous):

right. and when you simplified the numerator?

OpenStudy (anonymous):

well i got -10 ------ -5 which equals -2

OpenStudy (anonymous):

-10/-5 = 2

OpenStudy (anonymous):

what about this one? Given the functions k(x) = 5x − 8 and p(x) = x − 4, solve k[p(x)] and select the correct answer below. k[p(x)] = 5x − 12 k[p(x)] = 5x − 28 k[p(x)] = 5x2 − 12 k[p(x)] = 5x2 − 28

OpenStudy (anonymous):

@peachpi can you help?

OpenStudy (anonymous):

k(p(x)) means you substitute p(x) for the x in k(x)